OpenStudy (anonymous):

During one week, a man traveled 3-1/2, 1-1/4, 1-14/24, and 2-3/8 miles. The nest week he traveled 1/4, 3/8, 9/16, 3-1/16, 2-5/8, and 3-3/16 miles. How many more miles did he travel the second week than the first week?

OpenStudy (anonymous):

\[\text{Plus}\left[\frac{1}{4},\frac{3}{8},\frac{9}{16},3+\frac{1}{16},2+\frac{5}{8},3+\frac{3}{16}\right]-\text{Plus}\left[3+\frac{1}{2},1+\frac{1}{4},1+\frac{14}{24},2+\frac{3}{8}\right]= \] \[\frac{65}{48}=1.354 \]